OpenStudy (anonymous):

An ellipse has foci at (–1, 1) and (–1, –7). The major axis is 12 units long. Show work for all answers. -Determine the center of the ellipse. -Determine the vertices of the ellipse. -Find the equation of the ellipse.

OpenStudy (dumbcow):

only the "y-values" change with the foci , this tells you the major axis is vertical general equation of vertical ellipse \[\frac{(y-k)^{2}}{a^{2}} + \frac{(x-h)^{2}}{b^{2}} = 1\] where a^2 = b^2 + c^2 the distance between foci is 1-(-7) = 8 2c = 8 --> c = 4 --> c^2 = 16 given 2a = 12 --> a = 6 --> a^2 = 36 solve for b 36 = b^2 +16 b = sqrt(20) Now to find center, take midpoint of the foci --> (-1,-3) h = -1 , k = -3 \[\frac{(y+3)^{2}}{36}+\frac{(x+1)^{2}}{20} = 1\]